OpenStudy (anonymous):

Can someone help me with this special product? (x + 5)^2 I know I set it up as (x + 5) (x+5) and use FOIL. But so far I just have x^2, im just confused on combining x and 5 because it wouldnt be 5x right?

OpenStudy (anonymous):

yes. you should get \(5x\) twice, once from "outer" and once from "inner"

OpenStudy (anonymous):

btw it is not really "combining" it is "multiplying" because that is what you are doing when you multiply these two \(5x\) means \(5\times x\)

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

so would it be x^2 + 5x + 5x + 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now is when you "combine"

OpenStudy (anonymous):

but if I combined the 5x's to 10x^2?

OpenStudy (anonymous):

ok i made a mistake

OpenStudy (anonymous):

it is multiply, multiply, multiply, multiply not multiply, multiply, multiply, add

OpenStudy (anonymous):

\[x^2+5x+5x+5\times 5\] \[x^2+5x+5x+25\]

OpenStudy (anonymous):

now what is \(5x+5x\)? like $5 + $5

OpenStudy (anonymous):

oh okay! but then I combine the 5x's right?

OpenStudy (anonymous):

five dollars plus five dollars is ten dollars, and \(5x+5x=10x\)

OpenStudy (anonymous):

not \(10x^2\) when you add, you do not change the exponents

OpenStudy (anonymous):

okay, thank you very much!

OpenStudy (anonymous):

@chungerforever you got that? final answer should be \[x^2+10x+25\]

OpenStudy (anonymous):

with a little practice you will see that \((a+b)^2=a^2+2ab+b^2\) and will skip "foiling" so for example \[(x+6)^2=x^2+12x+36\]